OpenStudy (brucebaner):

Carrington is paid at the same rate each time he baby sits. The graph shows the rates he is paid for different times. Which equation represents the situation?

OpenStudy (brucebaner):

A. y=17x−5 B. y=17x+5 C. y=7x−5 D. y=7x+5

OpenStudy (brucebaner):

@kobeni-chan

OpenStudy (brucebaner):

@KamiBug

OpenStudy (brucebaner):

help me @KamiBug

OpenStudy (kobeni-chan):

You can tell that the y-intercept is 5, because 5 is where the line starts on the y-axis. To find the slope, take the coordinates on the graph and use the slope formula: (y-y)/(x-x)

OpenStudy (brucebaner):

so its A

OpenStudy (anonymous):

A. can someone give me a medal and then I will give you best answer

OpenStudy (brucebaner):

so its A @KamiBug and @kobeni-chan

OpenStudy (kobeni-chan):

yep @brucebaner

OpenStudy (kamibug):

@brucebaner The y-intercept is 5 and the slope is 7. So it's D. :)

OpenStudy (brucebaner):

thx

OpenStudy (kamibug):

We know the slope is 7 b/c we need to use the slope formula: 61-33 / 8-4 = 28/4 = 7

OpenStudy (brucebaner):

so its A @KamiBug

OpenStudy (kamibug):

It is D. A has the wrong slope.

OpenStudy (brucebaner):

D thx